# Parameter-Efficient Fine-Tuning for Continual Learning: A Neural Tangent Kernel Perspective

### v2

@@ 3 Preliminaries
$\Phi_{p_{\tau}}(x_{1},x_{2})=\left[J_{p_{\tau}}(f_{p_{\tau}}(x_{1}))\right]\left[J_{p_{\tau}}(f_{p_{\tau}}(x_{2}))\right]^{\top},$ (1)
where $J_{p_{\tau}}(f_{p_{\tau}}(x))$ denotes the Jacobian matrix of network $f_{\tau}$ with parameters optimized for task $\tau$, evaluated at input $x$. This function maps $D$-dimensional inputs to $O$-dimensional features, with $J_{p_{\tau}}(f_{p_{\tau}}(x))\in\mathbb{R}^{O\times P}$ and $\Phi_{p_{\tau}}(x_{1},x_{2})\in\mathbb{R}^{O\times O}$.
@@ 4 Theoretical Insights
$\footnotesize\begin{split}\mathcal{L}_{S}(f_{T}^{*}(X_{\tau}))\leq&\frac{1}{n_{\tau}}\bigg{[}\lambda^{2}\tilde{Y}_{\tau}^{\top}(\Phi_{\tau}(X_{\tau},X_{\tau})+\lambda I)^{-1}\tilde{Y}_{\tau}+\sum_{k=\tau+1}^{T}\tilde{Y}_{k}^{\top}\\ &\times(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\Phi_{k}(X_{\tau},X_{k})\\ &\times\Phi_{k}(X_{\tau},X_{k})^{\top}(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\tilde{Y}_{k}\bigg{]}_{\mathcal{D}_{\tau}},\end{split}$ (6)
^$\small\hat{\mathcal{R}}(\mathcal{F}_{T})\leq\left[\sum\limits_{\tau=1}^{T}\mathcal{O}(\sqrt{\frac{[\tilde{Y}_{\tau}^{\top}(\Phi_{\tau}(X,X)+\lambda I)^{-1}\tilde{Y}_{\tau}]}{n_{\tau}}})\right]_{\mathcal{D}_{\tau}}.$ (7)
@@ 5.2 Task-Level Feature Constraints
$[k_{1}(n),k_{2}(n)]=\begin{cases}[0,1]&\text{if }n=0\\ [\frac{1}{{n+1}}\cdot\frac{1}{k_{2}(n-1)},\frac{1}{{n+1}}]&\text{otherwise}\end{cases}$ (24)
@@ 6 Experiments
TABLE III: Comparative performance analysis in PEFT-CL using the ViT-Base16, fine-tuned on ImageNet-1K, as the foundational model. Bold segments indicate optimal results, while underlined segments denote suboptimal results.
@@
82.99 69.62
85.45 $\uparrow$+2.9670.45 $\uparrow$+1.19%
85.04 68.93
91.37 $\uparrow$+$7$uparrow+12.82%
86.51 71.93
89.50 $\uparrow$+3.4677.50 $\uparrow$+7.74%
89.47 76.13
92.01 $\uparrow$+2.8481.08 $\uparrow$+6.50%
93.32 $\uparrow$+4.3082.55 $\uparrow$+8.43%
92.39 $\uparrow$+3.2681.10 $\uparrow$+6.53%
92.39 $\uparrow$+3.2681.42 $\uparrow$+6.95%
93.47 $\uparrow$+4.4782.62 $\uparrow$+8.52%
93.72 $\uparrow$+4.7582.85 $\uparrow$+8.83%
@@ 6.1 Benchmark Comparison
Additionally, performance on auxiliary datasets including DomainNet, Oxford Pets, EuroSAT, PlantVillage, VTAB, and Kvasir, as detailed in Tables IV and V, highlights the generalization and adaptability of NTK-CL across diverse domains. On these datasets, NTK-CL not only consistently delivers superior accuracy metrics but also exhibits reduced variance in performance, emphasizing its stability. Notably, on Oxford Pets, NTK-CL achieves incremental accuracy improvements ranging from 1.8% to 2.1% and final accuracy enhancements of up to 4.6% compared to EASE [109]. On the Kvasir dataset, NTK-CL outperforms competing methods, achieving the highest incremental accuracy improvements ranging from 6.7% to 9.0% and the highest final accuracy improvements ranging from 19.3% to 21.1%, showcasing its significant potential for medical applications. Across other datasets, NTK-CL consistently ranks as the best or the second-best, further affirming the method’s efficacy and versatility.
@@ 6.2 Ablation Study
To rigorously align theoretical constructs with empirical evidence, an extensive series of ablation studies are conducted utilizing the CIFAR100 and ImageNet-R datasets. All experiments adhere to standardized conditions: a fixed random seed of 0, consistent task segmentation, and the utilization of pre-trained ViT-B/16-IN21K weight to ensure model consistency. The ablation studies examine configurations involving the Subnetwork-1 Adaptation Module (S1), Subnetwork-2 Adaptation Module (S2), Hybrid Adaptation Module (Hybrid), Knowledge Retention (KR) mechanism, Task-Feature Dissimilarity Loss (Dis), Orthogonality Loss (Orth), and Regularization Loss (Reg). The average accuracy ($\bar{A}$) across tasks is evaluated to quantitatively assess the contributions of each component to the model’s overall performance, which is displayed in Table VI.
Preliminary analyses reveal that the Hybrid module surpasses the standalone S1 and S2 modules by synergistically combining their strengths. This synergy, when optimized jointly, significantly boosts performance, highlighting the benefits of increased sample diversity in promoting knowledge transfer and retention within the PEFT-CL scenario. The efficacy of the KR module, in conjunction with different adaptation configurations, is systematically investigated. Results indicate that integrating the KR module markedly improves $\bar{A}$ across all adaptation modules. Notably, the KR module elevates the $\bar{A}$ from 82.99% to 85.45% on CIFAR100 and from 69.62% to 70.45% on ImageNet-R for the S1 configuration. For S2 module, the improvement is more pronounced, increasing from 85.04% to 91.37% on CIFAR100 and from 68.93% to 77.77% on ImageNet-R. The Hybrid module also sees a significant boost, with $\bar{A}$ rising from 86.51% to 89.50% on CIFAR100 and from 71.93% to 77.50% on ImageNet-R. Combining all three adaptation modules with the KR module achieves the highest $\bar{A}$ of 92.01% on CIFAR100 and 81.08% on ImageNet-R, underscoring the KR module’s pivotal role in enhancing knowledge retention and model generalization. The introduction of task-feature dissimilarity loss (Dis) further enhances performance, achieving $\bar{A}$ of 93.32% on CIFAR100 and 82.55% on ImageNet-R, representing improvements of 4.30% and 8.43%, respectively. Incorporating orthogonality loss (Orth) alongside the adaptation modules and KR module yields an $\bar{A}$ of 92.39% on CIFAR100 and 81.10% on ImageNet-R, indicating gains of 3.26% and 6.53%. Adding regularization loss (Reg) to this configuration further refines performance, achieving an $\bar{A}$ of 92.39% on CIFAR100 and 81.42% on ImageNet-R, with improvements of 3.26% and 6.95%. Integrating all components and strategies culminates in peak $\bar{A}$ values of 93.72% on CIFAR100 and 82.85% on ImageNet-R, marking cumulative enhancements of 4.75% and 8.83%. In summary, these ablation studies validate our theoretical hypotheses, confirming the effectiveness of each design choice.
@@
TABLE VII: Comparison of feature size expansion methods and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Combinations Type Incremental Top-1 Accuracy
@@
Original + Mixed-up [105] 98.50 93.40 93.12 92.80 91.90 89.40 88.23 87.95 87.24 87.09
Original + PuzzleMix [41] 98.50 93.75 93.07 92.92 91.58 89.39 88.48 88.16 87.19 87.08
Original + AutoAug [16] 98.20 94.15 93.28 93.13 91.72 90.19 88.82 88.48 87.50 87.24
Original + RandAug [17] 98.20 94.05 93.42 93.10 92.02 90.41 89.80 89.18 88.09 87.99
IA3 [53] + S1 96.00 65.04 63.00 61.35 60.77 60.65 60.32 59.56 58.35 53.06
IA3 [53] + S2 98.50 96.35 95.07 94.20 93.54 91.21 91.18 89.94 89.09 88.59
Compacter [58] + S1 97.90 95.85 94.73 94.30 93.48 91.63 91.23 90.89 90.06 89.48
Compacter [58] + S2 98.60 96.75 95.37 94.80 94.30 92.04 92.03 91.26 90.59 90.03
Side-Tuning [106] + S1 95.80 93.55 91.70 89.48 88.36 85.89 85.82 85.06 84.02 83.18
Side-Tuning [106] + S2 98.50 96.55 95.73 94.10 93.62 91.33 91.20 90.46 89.92 89.11
Side-Tuning [106] + Compacter [58] 98.00 96.10 95.17 94.78 93.96 91.74 91.28 90.85 90.08 89.25
S1 + S2 (Ours) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE VIII: Comparison of feature fusion methods and their impact on the evolution of incremental top-1 accuracy. $Q$, $K$, and $V$ denote the query, key, and value in our feature fusion method. The structures of the various fusion methods are illustrated in Appendix J. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Fusion Type Incremental Top-1 Accuracy
@@
MLP 98.90 96.35 94.70 93.98 93.88 91.22 90.90 89.75 88.83 88.49
VAE 98.60 96.25 95.23 94.62 93.92 91.49 91.42 90.52 89.81 89.00
RNN 98.80 96.45 95.07 94.88 94.00 91.63 91.29 90.45 89.97 89.17
Mamba 98.30 96.75 95.50 94.68 93.84 91.63 91.24 90.20 89.86 89.12
S1($K/V$) + S2 ($Q$) 98.70 96.55 95.57 94.98 94.10 91.73 91.54 90.91 90.48 90.05
S1($Q$) + S2 ($K/V$) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE IX: Comparison of knowledge inheritance methods and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Knowledge Inheritance Type Incremental Top-1 Accuracy
@@
MoCo-v3 [13] 98.70 95.90 93.17 91.32 90.50 87.63 87.31 87.01 86.21 85.36
LAE [22] 98.70 96.05 93.37 91.48 90.44 87.68 87.17 86.92 86.12 85.40
EASE [109] 98.70 96.65 91.77 91.45 91.36 91.15 90.89 90.49 89.36 88.23
Adaptive EMA (Ours) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE X: Comparison of orthogonality losses and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Orthogonality Loss Type Incremental Top-1 Accuracy
@@
$\mathcal{L}_{\text{Standard}}$ Class-level 98.70 96.60 95.80 94.62 94.42 92.08 92.01 91.28 90.43 89.98
$\mathcal{L}_{\text{Schmidt}}$ Class-level 98.70 96.50 95.63 94.62 94.04 91.68 91.46 90.60 89.67 89.10
$\mathcal{L}_{\text{orth}}$ (Ours) Task-level 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24
TABLE XI: Comparison of regularization methods and their impact on the evolution of incremental top-1 accuracy. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Regularization Type Incremental Top-1 Accuracy
@@
L1 97.60 96.00 94.97 94.02 93.68 91.47 91.17 90.41 89.70 88.76
Spectral 98.60 96.60 95.70 94.70 94.34 92.04 92.00 91.12 90.57 89.06
HiDe-Prompt [82] 98.70 96.55 95.53 94.70 94.22 92.02 91.89 91.15 90.72 90.11
L2 (Ours) 98.70 96.65 95.70 94.85 94.36 92.28 92.07 91.44 90.90 90.24

### v3

@@ 3 Preliminaries
$\Phi_{p_{\tau}}(x_{1},x_{2})=\left[J_{p_{\tau}}(f_{\tau}(x_{1}))\right]\left[J_{p_{\tau}}(f_{\tau}(x_{2}))\right]^{\top},$ (1)
where $J_{p_{\tau}}(f_{\tau}(x))$ denotes the Jacobian matrix of network $f_{\tau}$ with parameters optimized for task $\tau$, evaluated at input $x$. This function maps $D$-dimensional inputs to $O$-dimensional features, with $J_{p_{\tau}}(f_{\tau}(x))\in\mathbb{R}^{O\times P}$ and $\Phi_{p_{\tau}}(x_{1},x_{2})\in\mathbb{R}^{O\times O}$.
@@ 4 Theoretical Insights
$\footnotesize\begin{split}L_{S}(f_{T}^{*}(X_{\tau}))\leq&\frac{1}{n_{\tau}}\bigg[\lambda^{2}\tilde{Y}_{\tau}^{\top}(\Phi_{\tau}(X_{\tau},X_{\tau})+\lambda I)^{-1}\tilde{Y}_{\tau}+\sum_{k=\tau+1}^{T}\tilde{Y}_{k}^{\top}\\ &\times(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\Phi_{k}(X_{\tau},X_{k})\\ &\times\Phi_{k}(X_{\tau},X_{k})^{\top}(\Phi_{k}(X_{k},X_{k})+\lambda I)^{-1}\tilde{Y}_{k}\bigg]_{\mathcal{D}_{\tau}},\end{split}$ (6)
$\small\hat{\mathcal{R}}(\mathcal{F}_{T})\leq\left[\sum\limits_{\tau=1}^{T}\mathcal{O}(\sqrt{\frac{[\tilde{Y}_{\tau}^{\top}(\Phi_{\tau}(X,X)+\lambda I)^{-1}\tilde{Y}_{\tau}]}{n_{\tau}}})\right]_{\mathcal{D}_{\tau}}.$ (7)
@@ 5.2 Task-Level Feature Constraints
[…]
@@ 6 Experiments
TABLE III: Comparative performance analysis in PEFT-CL using ViT-Base16, fine-tuned on ImageNet-1K, as the foundational model. Bold segments indicate optimal results, while underlined segments denote suboptimal results.
@@
83.22 $\pm$ 0.41 67.78 $\pm$ 2.99
84.61 $\pm$ 1.69 69.89 $\pm$ 2.35
85.14 $\pm$ 0.35 68.84 $\pm$ 0.33
91.29 $\pm$ 0.70 77.63 $\pm$ 0.51
86.75 $\pm$ 0.32 71.90 $\pm$ 0.29
88.92 $\pm$ 0.60 77.39 $\pm$ 0.58
89.18 $\pm$ 0.35 75.94 $\pm$ 0.27
91.89 $\pm$ 0.36 80.91 $\pm$ 0.31
92.83 $\pm$ 0.56 82.12 $\pm$ 0.59
92.39 $\pm$ 0.36 81.26 $\pm$ 0.41
92.16 $\pm$ 0.36 81.41 $\pm$ 0.41
93.06 $\pm$ 0.55 82.50 $\pm$ 0.61
93.76 $\pm$ 0.35 82.77 $\pm$ 0.66
@@ 6.1 Benchmark Comparison
Additionally, performance on auxiliary datasets including DomainNet, Oxford Pets, EuroSAT, PlantVillage, VTAB, and Kvasir, as detailed in Tables IV and V, highlights the generalization and adaptability of NTK-CL across diverse domains. On these datasets, NTK-CL not only consistently delivers superior accuracy metrics but also exhibits reduced standard deviation in performance, emphasizing its stability. Notably, on Oxford Pets, NTK-CL achieves incremental accuracy improvements ranging from 1.8% to 2.1% and final accuracy enhancements of up to 4.6% compared to EASE [109]. On the Kvasir dataset, NTK-CL outperforms competing methods, achieving the highest incremental accuracy improvements ranging from 6.7% to 9.0% and the highest final accuracy improvements ranging from 19.3% to 21.1%, showcasing its significant potential for medical applications. Across other datasets, NTK-CL consistently ranks as the best or the second-best, further affirming the method’s efficacy and versatility.
@@ 6.2 Ablation Study
To systematically evaluate the contribution of each component in the proposed framework, we conduct extensive ablation studies on CIFAR100 and ImageNet-R using a frozen ViT-B/16-IN21K backbone. For fair comparison, all experiments are conducted under identical settings, including fixed random seeds (0-4), consistent task partitions, and the same pre-trained backbone initialization. The ablation studies examine configurations involving the Subnetwork-1 Adaptation Module (S1), Subnetwork-2 Adaptation Module (S2), Hybrid Adaptation Module (Hybrid) 666“Hybrid Only” refers to the training objective rather than the network structure. In this setting, both Subnetwork-1 (S1) and Subnetwork-2 (S2) are retained in the model and remain fully trainable, producing two adapted feature streams. The Hybrid Adaptation Module performs cross-attention based fusion between these two streams to obtain $E_{CLS}^{HAE}$, rather than operating on a single backbone output with $Q=K=V$. The term “Hybrid Only” indicates that the model is supervised solely using $CE(E_{CLS}^{HAE},y)$ in Eq. 23, while the branch-specific auxiliary losses on S1 and S2 are disabled., Knowledge Retention (KR) mechanism, Task-Feature Dissimilarity Loss (Dis), Orthogonality Loss (Orth), and Regularization Loss (Reg). The average accuracy ($\bar{A}$) across tasks is evaluated to quantitatively assess the contributions of each component.
The results in Table VI indicate that different adaptation strategies consistently improve performance over the frozen backbone. Among them, the Hybrid adaptation module achieves higher average accuracy than the standalone Subnetwork-1 (S1) and Subnetwork-2 (S2) configurations on both datasets, suggesting that jointly leveraging multiple adaptation pathways is more effective for continual representation learning. The introduction of the Knowledge Retention (KR) mechanism further improves performance across all adaptation configurations. In particular, for the S2-based setting, incorporating KR leads to a notable increase in average accuracy on both CIFAR100 and ImageNet-R, while for the S1 and Hybrid configurations, KR provides consistent performance gains with comparable variability across runs. These results indicate that KR effectively enhances knowledge preservation under different adaptation strategies. We further analyze the effect of auxiliary constraints, including the task-feature dissimilarity loss, orthogonality loss, and regularization loss. When combined with the adaptation modules and KR, these components yield additional improvements in average accuracy on both datasets. The performance variations reported in Table VI remain within a comparable range across different configurations, indicating that the observed gains are stable and reproducible. Finally, integrating all adaptation modules, the KR mechanism, and auxiliary losses achieves the best overall performance on both CIFAR100 and ImageNet-R. These results demonstrate that each component contributes to the final performance, with the Hybrid adaptation and Knowledge Retention mechanisms playing a central role in the proposed parameter-efficient continual learning framework.
@@
TABLE VII: Comparison of feature size expansion methods on CIFAR100 and their impact on the evolution of incremental top-1 accuracy (%). The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Combinations Type Incremental Top-1 Accuracy (%)
@@
Original + Mixed-up [105] 98.50 $\pm$ 0.98 94.70 $\pm$ 1.64 93.21 $\pm$ 1.15 91.99 $\pm$ 0.95 90.19 $\pm$ 2.06 90.07 $\pm$ 1.30 88.52 $\pm$ 0.97 87.81 $\pm$ 0.33 86.92 $\pm$ 0.74 86.91 $\pm$ 0.62
Original + PuzzleMix [41] 98.44 $\pm$ 0.90 94.98 $\pm$ 1.42 93.53 $\pm$ 0.77 93.26 $\pm$ 0.47 90.26 $\pm$ 1.74 90.04 $\pm$ 1.11 88.65 $\pm$ 0.92 88.41 $\pm$ 0.65 86.72 $\pm$ 0.79 86.60 $\pm$ 0.77
Original + AutoAug [17] 98.50 $\pm$ 0.85 95.14 $\pm$ 1.47 93.54 $\pm$ 1.27 92.39 $\pm$ 0.96 91.94 $\pm$ 1.96 91.07 $\pm$ 1.27 89.39 $\pm$ 1.14 88.82 $\pm$ 0.85 87.98 $\pm$ 0.96 87.06 $\pm$ 0.52
Original + RandAug [17] 98.54 $\pm$ 0.91 95.65 $\pm$ 1.87 93.38 $\pm$ 1.17 92.81 $\pm$ 0.93 90.17 $\pm$ 2.19 90.05 $\pm$ 1.28 89.54 $\pm$ 0.88 88.67 $\pm$ 0.53 87.94 $\pm$ 0.67 87.89 $\pm$ 0.63
IA3 [53] + S1 96.72 $\pm$ 1.31 65.77 $\pm$ 3.20 64.20 $\pm$ 4.63 60.83 $\pm$ 2.95 59.70 $\pm$ 3.88 59.95 $\pm$ 5.11 58.74 $\pm$ 4.51 58.09 $\pm$ 2.85 57.20 $\pm$ 2.64 55.38 $\pm$ 3.15
IA3 [53] + S2 98.48 $\pm$ 0.85 97.05 $\pm$ 0.79 95.26 $\pm$ 0.77 94.22 $\pm$ 0.40 93.02 $\pm$ 1.13 91.61 $\pm$ 0.98 91.21 $\pm$ 0.48 90.46 $\pm$ 0.80 89.41 $\pm$ 0.50 88.96 $\pm$ 0.56
Compacter [58] + S1 98.24 $\pm$ 1.16 96.55 $\pm$ 0.99 94.69 $\pm$ 0.93 93.80 $\pm$ 0.59 92.79 $\pm$ 0.86 92.32 $\pm$ 0.94 91.06 $\pm$ 0.65 90.36 $\pm$ 0.54 89.65 $\pm$ 0.58 89.18 $\pm$ 0.48
Compacter [58] + S2 98.54 $\pm$ 0.89 97.35 $\pm$ 0.65 95.65 $\pm$ 0.70 94.57 $\pm$ 0.37 93.61 $\pm$ 0.87 92.17 $\pm$ 0.99 91.92 $\pm$ 0.44 91.29 $\pm$ 0.38 90.50 $\pm$ 0.22 90.10 $\pm$ 0.18
Side-Tuning [106] + S1 96.06 $\pm$ 1.30 93.45 $\pm$ 1.41 90.47 $\pm$ 1.90 87.93 $\pm$ 1.62 86.49 $\pm$ 2.22 85.95 $\pm$ 1.01 84.52 $\pm$ 1.80 83.89 $\pm$ 1.76 82.98 $\pm$ 1.69 82.58 $\pm$ 0.81
Side-Tuning [106] + S2 98.48 $\pm$ 0.99 97.42 $\pm$ 0.90 95.71 $\pm$ 0.67 94.20 $\pm$ 0.31 93.52 $\pm$ 0.75 92.23 $\pm$ 0.93 91.03 $\pm$ 0.28 90.42 $\pm$ 0.34 89.58 $\pm$ 0.39 89.22 $\pm$ 0.19
Side-Tuning [106] + Compacter [58] 98.38 $\pm$ 1.07 96.84 $\pm$ 0.81 94.87 $\pm$ 0.87 93.87 $\pm$ 0.98 92.79 $\pm$ 1.42 92.41 $\pm$ 0.72 91.31 $\pm$ 0.59 90.66 $\pm$ 0.33 89.80 $\pm$ 0.47 89.41 $\pm$ 0.31
S1 + S2 (Ours) 98.52 $\pm$ 0.98 97.33 $\pm$ 0.71 95.69 $\pm$ 0.76 94.72 $\pm$ 0.28 93.65 $\pm$ 0.88 93.25 $\pm$ 0.97 91.99 $\pm$ 0.65 91.44 $\pm$ 0.40 90.71 $\pm$ 0.24 90.27 $\pm$ 0.20
TABLE VIII: Comparison of feature fusion methods on CIFAR100 and their impact on incremental top-1 accuracy (%). $Q$, $K$, and $V$ are the query, key, and value in our feature fusion method. The structures of the various fusion methods are illustrated in Appendix J. The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Fusion Type Incremental Top-1 Accuracy (%)
@@
MLP 98.32 $\pm$ 1.20 96.83 $\pm$ 1.05 95.13 $\pm$ 0.70 93.63 $\pm$ 0.73 92.61 $\pm$ 1.49 92.03 $\pm$ 0.82 90.31 $\pm$ 0.97 89.46 $\pm$ 0.40 88.46 $\pm$ 0.62 88.25 $\pm$ 0.46
VAE 98.16 $\pm$ 1.40 96.67 $\pm$ 0.84 94.43 $\pm$ 0.87 93.49 $\pm$ 1.74 92.18 $\pm$ 1.80 91.83 $\pm$ 0.88 90.38 $\pm$ 1.10 89.77 $\pm$ 0.79 88.92 $\pm$ 1.73 88.57 $\pm$ 0.46
RNN 98.10 $\pm$ 1.38 96.54 $\pm$ 0.70 94.83 $\pm$ 0.42 94.66 $\pm$ 0.69 92.33 $\pm$ 1.70 92.05 $\pm$ 0.92 90.75 $\pm$ 0.78 90.08 $\pm$ 0.38 89.31 $\pm$ 0.74 88.96 $\pm$ 0.42
Mamba 98.00 $\pm$ 1.14 96.86 $\pm$ 0.90 95.30 $\pm$ 0.99 94.30 $\pm$ 0.57 91.80 $\pm$ 2.48 91.16 $\pm$ 1.87 89.83 $\pm$ 1.64 89.04 $\pm$ 1.39 89.17 $\pm$ 1.04 87.81 $\pm$ 1.39
S1($K/V$) + S2 ($Q$) 97.72 $\pm$ 1.43 96.14 $\pm$ 1.26 94.31 $\pm$ 1.33 93.25 $\pm$ 1.90 92.51 $\pm$ 2.97 90.57 $\pm$ 1.78 90.46 $\pm$ 1.40 90.02 $\pm$ 1.00 89.51 $\pm$ 1.10 89.08 $\pm$ 1.19
S1($Q$) + S2 ($K/V$) 98.52 $\pm$ 0.98 97.33 $\pm$ 0.71 95.69 $\pm$ 0.76 94.72 $\pm$ 0.28 93.65 $\pm$ 0.88 93.25 $\pm$ 0.97 91.99 $\pm$ 0.65 91.44 $\pm$ 0.40 90.71 $\pm$ 0.24 90.27 $\pm$ 0.20
TABLE IX: Comparison of knowledge inheritance methods on CIFAR100 and their impact on incremental top-1 accuracy (%). The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Knowledge Inheritance Type Incremental Top-1 Accuracy (%)
@@
MoCo-v3 [13] 98.08 $\pm$ 1.52 96.12 $\pm$ 1.34 92.93 $\pm$ 1.69 91.11 $\pm$ 1.96 89.62 $\pm$ 3.05 88.83 $\pm$ 1.62 88.15 $\pm$ 1.89 87.33 $\pm$ 1.49 86.82 $\pm$ 1.17 86.70 $\pm$ 1.38
LAE [22] 98.00 $\pm$ 1.49 94.91 $\pm$ 1.35 92.48 $\pm$ 1.64 89.93 $\pm$ 2.52 87.29 $\pm$ 3.62 86.48 $\pm$ 3.51 85.21 $\pm$ 3.07 84.70 $\pm$ 2.57 83.94 $\pm$ 2.30 83.56 $\pm$ 2.61
EASE [109] 98.24 $\pm$ 1.18 96.53 $\pm$ 1.08 94.24 $\pm$ 2.80 92.19 $\pm$ 0.75 90.95 $\pm$ 1.89 90.29 $\pm$ 1.04 89.69 $\pm$ 1.57 87.79 $\pm$ 2.87 87.19 $\pm$ 2.30 87.14 $\pm$ 1.74
Adaptive EMA (Ours) 98.52 $\pm$ 0.98 97.33 $\pm$ 0.71 95.69 $\pm$ 0.76 94.72 $\pm$ 0.28 93.65 $\pm$ 0.88 93.25 $\pm$ 0.97 91.99 $\pm$ 0.65 91.44 $\pm$ 0.40 90.71 $\pm$ 0.24 90.27 $\pm$ 0.20
TABLE X: Comparison of orthogonality losses on CIFAR100 and their impact on incremental top-1 accuracy (%). The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Orthogonality Loss Type Incremental Top-1 Accuracy (%)
@@
$\mathcal{L}_{\text{Standard}}$ Class-level 98.46 $\pm$ 0.99 97.33 $\pm$ 0.73 95.70 $\pm$ 0.78 94.53 $\pm$ 0.47 93.47 $\pm$ 0.97 92.10 $\pm$ 0.75 91.71 $\pm$ 0.54 91.16 $\pm$ 0.33 90.30 $\pm$ 0.21 89.90 $\pm$ 0.21
$\mathcal{L}_{\text{Schmidt}}$ Class-level 98.46 $\pm$ 0.91 97.30 $\pm$ 0.88 95.56 $\pm$ 0.77 94.20 $\pm$ 0.62 92.89 $\pm$ 1.29 92.61 $\pm$ 0.99 91.02 $\pm$ 0.81 90.22 $\pm$ 0.60 89.26 $\pm$ 0.52 88.76 $\pm$ 0.51
$\mathcal{L}_{\text{orth}}$ (Ours) Task-level 98.52 $\pm$ 0.98 97.33 $\pm$ 0.71 95.69 $\pm$ 0.76 94.72 $\pm$ 0.28 93.65 $\pm$ 0.88 93.25 $\pm$ 0.97 91.99 $\pm$ 0.65 91.44 $\pm$ 0.40 90.71 $\pm$ 0.24 90.27 $\pm$ 0.20
TABLE XI: Comparison of regularization methods on CIFAR100 and their impact on incremental top-1 accuracy (%). The bold segments denote optimal results, and the underlined segments indicate suboptimal outcomes.
Regularization Type Incremental Top-1 Accuracy (%)
@@
L1 97.34 $\pm$ 1.44 95.98 $\pm$ 1.10 94.41 $\pm$ 0.59 93.55 $\pm$ 0.97 92.40 $\pm$ 1.45 91.80 $\pm$ 0.60 90.60 $\pm$ 0.69 89.87 $\pm$ 0.55 89.16 $\pm$ 0.73 88.67 $\pm$ 0.25
Spectral 98.54 $\pm$ 0.89 97.27 $\pm$ 0.75 95.75 $\pm$ 0.82 94.67 $\pm$ 0.29 93.56 $\pm$ 0.90 93.16 $\pm$ 1.80 91.90 $\pm$ 0.61 91.27 $\pm$ 0.32 90.53 $\pm$ 0.16 89.14 $\pm$ 0.15
HiDe-Prompt [82] 98.56 $\pm$ 0.90 97.12 $\pm$ 0.63 95.62 $\pm$ 0.74 94.56 $\pm$ 0.38 93.44 $\pm$ 0.84 92.58 $\pm$ 0.76 91.82 $\pm$ 0.55 91.23 $\pm$ 0.45 90.50 $\pm$ 0.35 90.04 $\pm$ 0.27
L2 (Ours) 98.52 $\pm$ 0.98 97.33 $\pm$ 0.71 95.69 $\pm$ 0.76 94.72 $\pm$ 0.28 93.65 $\pm$ 0.88 93.25 $\pm$ 0.97 91.99 $\pm$ 0.65 91.44 $\pm$ 0.40 90.71 $\pm$ 0.24 90.27 $\pm$ 0.20